OpenStudy (anonymous):

the square of a number plus the number is 6. find the number

OpenStudy (anonymous):

x^2+x=6 ===> x^2+x-6=0 ===>( x-3)(x+2)=0 ==> x=3 or x=-2

OpenStudy (anonymous):

The answer is 9. 9/3+3=6

OpenStudy (anonymous):

correction: x^2+x=6 ===> x^2+x-6=0 ===>( x+3)(x-2)=0 ==> x=-3 or x=2